Given the following Revenue and Cost Functions:
Revenue = -8.0X2 + 200.0X
Cost = 45.0X + 20.0
Find the Highest Breakeven Point.

Answer by stanbon(75887)

Given the following Revenue and Cost Functions:
Revenue = -8.0X2 + 200.0X
Cost = 45.0X + 20.0
Find the Highest Breakeven Point.
--------------
Breakeven occurs when Revenue = Cost
-8.0x^2 + 200x = 45x + 20
-------
8x^2 -155x + 20 = 0
-------
Use the Quadratic Formula to get::
x = 15.25